OpenStudy (anonymous):

Find local min and max and the points of inflection for f(x)= 3(cos(x))^2 - 6sin(x)

OpenStudy (anonymous):

Do you know how to do the chain rule with respect to trig functions?

OpenStudy (anonymous):

Example \[ \bigg(\cos^4(x)\bigg)^{\prime}=4\cos^3(x)\]

OpenStudy (anonymous):

oh and generally \[\cos^2(x) = (\cos(x))^2\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

let me write it correctly \[f'(x)=-6\cos(x)\sin(x)-6\cos(x)]\] that's better

myininaya (myininaya):

\[\(cos^2(x))'=2\cos(x)(-\sin(x))=-2\cos(x)\sin(x)\]

myininaya (myininaya):

@freewilly922 your example is a little off

OpenStudy (anonymous):

mine was way off!

myininaya (myininaya):

Corrected example from above: \[(\cos^4(x))'=4\cos^3(x)(-\sin(x))=-4\cos^3(x)\sin(x)\]

OpenStudy (anonymous):

i just ate a mini kitkat